For the production function Q = min{4L,3K}, what bundle of inputs is on the same isoquant as (L=4,K=4)?

(a.) (L=2,K=3)

(b.) (L=4, K=1)

(c.) (L=2, K=1)

(d.) (L=6,K=2)
Answer #1

Q = min{4L,3K} and we have to find the bundle of inputs which is on the same isoquant with (L=4,K=4).

The input bundle which produces the same output(Q) as (L=4,K=4) will be on the same isoquant with it.

First of all we have to find out the output produced by the input bundle (L=4,K=4) which is

Q = min{4L,3K}= min{4*4, 3*4} = min{16, 12} = 12

Now we will check each option particularly whether it is producing the same output or not.

a) (L=2, K=3)

Q=min{4L,3K}= min{4*2, 3*3}= min{8,9}= 8

b) (L=4, K=1)

Q=min{4L,3K}= min{4*4, 3*1}= min{16, 3}= 3

c) (L=2, K=1)

Q= min{4L,3K}= min{4*2, 3*1}= min{8, 3}= 3

d) (L=6, K=2)

Q= min{4L,3K}= min{4*6, 3*2}= min{24, 6}= 6

As we can see none of the options are producing 12 units of output. According to this question none of the option is correct.